# Monopole-vortex continuity of ${\mathcal{N}}=1$ super Yang-Mills theory on $\mathbb{R}^{2}\times S^{1}\times S^{1}$ with ’t Hooft twist

### v2

@@ Abstract
We study ${\mathcal{N}}=1$ $SU(N)$ super Yang-Mills (SYM) theory on $\mathbb{R}^{2}\times(S^{1})_{3}\times(S^{1})_{4}$ with the ’t Hooft twist. The theory becomes weakly coupled if the length $L_{4}$ of $(S^{1})_{4}$ is sufficiently small, $NL_{4}\Lambda\ll 1$. We explore the nonperturbative dynamics at the weak-coupling regime by changing the size of $L_{3}$ and uncover how $3$d monopole/bion-based effective theory for $L_{3}\gg L_{4}$ is related to the $2$d vortex-based theory for $L_{3}\approx L_{4}$. The highlights of our results are (1) the smooth “weak-weak” continuity of the vacuum structure and gluino condensate during the $3$d-$2$d dimensional reduction, (2) the crossover of Wilson loop behavior from the area law in $3$d to the perimeter law in $2$d via a “double-string” picture, (3) the role of mass deformation in breaking discrete chiral symmetry and restoring the area law in $2$d, and (4) the microscopic investigation of bions during the reduction from $3$d to $2$d and the cancellation of the vacuum energy due to the hidden topological angle. We also discuss the generalization of our results for (1)–(3) from $\mathcal{N}=1$ SYM to QCD with adjoint quarks.
@@ 1 Introduction
Double string picture and the transition from 3d area law to 2d perimeter law: In Section 4, we investigate the behavior of the Wilson loop during the dimensional transition from ${\mathbb{R}}^{3}\times S^{1}$ to ${\mathbb{R}}^{2}\times T^{2}$ with ’t Hooft flux. The 4d mixed anomaly between the center and chiral symmetry produces the 2d mixed anomaly between the 1-form center and chiral symmetry under the presence of ’t Hooft flux, which requires the 2d perimeter law of the Wilson loop. We show that the perimeter law of the 2d center-vortex semiclassics can be understood from the 3d monopole through the “double-string picture”. With the double-string picture, the Wilson loop creates the domain wall for the broken chiral symmetry that wraps along the $S^{1}$ direction, and the Wilson loop serves as the junction of the domain wall. With the mass deformation, Wilson loops obey the area law due to the absence of the discrete chiral symmetry.
@@ 2 Review on $\mathbb{R}^{3}\times S^{1}$ semiclassics and $\mathbb{R}^{2}\times T^{2}$ semiclassics
To confirm these claims semiclassically, a convenient way to describe the vacuum structure is to compactify $\mathbb{R}^{2}\times(T^{2})_{34}$ to ×$\mathbb{R}\times(S^{1})_{2}\times(T^{2})_{34}$ with the size $L_{2}$, and we set $L=L_{3,4}\ll L_{2}\ll\Lambda^{-1}$. We then take the Hamiltonian viewpoint with regarding $\mathbb{R}$ as the time direction for the quantization Tanizaki:2022ngt ; Yamazaki:2017ulc ; Cox:2021vsa . As $L=L_{3,4}$ is the shortest length scale, we first minimize along those directions, and the ’t Hooft twist $n_{34}=1$ sets $P_{3}=S$, $P_{4}=C$ as we have already seen. Next, $P_{2}$ must commute with both $P_{3},P_{4}$ to minimize the classical action along the $2$-$3$ and $2$-$4$ directions, and then we obtain $N$ classical vacua labelled by $P_{2}=\mathrm{e}^{\frac{2\pi\mathrm{i}m}{N}}1_{N\times N}$ with $m=0,1,\ldots,N-1$. Let us call this state |$\ket{m}$:
@@
  • It is quite counter-intuitive that the Wilson loop in $2$d obeys the perimeter law whereas the vacuum structure is smoothly connected to the $4$d/$3$d confinement vacua. What is the microscopic dynamics realizing the crossover from the 3d area law to the 2d perimeter law?
@@ 4.1.2 Wilson loop in ${\mathbb{R}}^{2}\times S^{1}$: perimeter law from double string picture
where $T_{\mathrm{kink}}$ is the kink tension, and $T_{\mathrm{conf}}$ is the string tension. If the size of $C$ is not so large compared with $L_{3}$, the standard area law appears. When the size of $C$ becomes large enough compared to $L_{3}$, the domain wall extended along $(S^{1})_{3}$ dominates and we obtain the perimeter law. The crossover scale between these behaviors can be estimated as $T_{\mathrm{conf}}\,\mathrm{Area}(C_{*})=T_{\mathrm{kink}}\,L_{3}\,\mathrm{Length}(C_{*})$; when we use the double-string scenario with $T_{\mathrm{conf}}\approx 2T_{\mathrm{kink}}$, the crossover size becomes $L_{*}\approx 2L_{3}$ for a square loop of size $L\times L$. This is the microscopic explanation for why the Wilson loops in the $2$d center-vortex semiclassics shows the perimeter law and generates the spontaneously-broken chiral transformation.

### v3

@@ Abstract
We study ${\mathcal{N}}=1$ $SU(N)$ super Yang-Mills (SYM) theory on $\mathbb{R}^{2}\times(S^{1})_{3}\times(S^{1})_{4}$ with the ’t Hooft twist. The theory becomes weakly coupled if the length $L_{4}$ of $(S^{1})_{4}$ is sufficiently small, $NL_{4}\Lambda\ll 1$. We explore the nonperturbative dynamics at the weak-coupling regime by changing the size of $L_{3}$ and uncover how $3$d monopole/bion-based effective theory for $L_{3}\gg L_{4}$ is related to the $2$d vortex-based theory for $L_{3}\approx L_{4}$. The highlights of our results are (1) the smooth “weak-weak” continuity of the vacuum structure and gluino condensate during the $3$d-$2$d dimensional reduction, (2) the switching of Wilson loop behavior from the area law in $3$d to the perimeter law in $2$d via a “double-string” picture, (3) the role of mass deformation in breaking discrete chiral symmetry and restoring the area law in $2$d, and (4) the microscopic investigation of bions during the reduction from $3$d to $2$d and the cancellation of the vacuum energy due to the hidden topological angle. We also discuss the generalization of our results for (1)–(3) from $\mathcal{N}=1$ SYM to QCD with adjoint quarks.
@@ 1 Introduction
Double string picture and the switching from 3d area law to 2d perimeter law: In Section 4, we investigate the behavior of the Wilson loop during the dimensional transition from ${\mathbb{R}}^{3}\times S^{1}$ to ${\mathbb{R}}^{2}\times T^{2}$ with ’t Hooft flux. The 4d mixed anomaly between the center and chiral symmetry produces the 2d mixed anomaly between the 1-form center and chiral symmetry under the presence of ’t Hooft flux, which requires the 2d perimeter law of the Wilson loop. We show that the perimeter law of the 2d center-vortex semiclassics can be understood from the 3d monopole through the “double-string picture”. With the double-string picture, the Wilson loop creates the domain wall for the broken chiral symmetry that wraps along the $S^{1}$ direction, and the Wilson loop serves as the junction of the domain wall. With the mass deformation, Wilson loops obey the area law due to the absence of the discrete chiral symmetry.
@@ 2 Review on $\mathbb{R}^{3}\times S^{1}$ semiclassics and $\mathbb{R}^{2}\times T^{2}$ semiclassics
To confirm these claims semiclassically, a convenient way to describe the vacuum structure is to compactify $\mathbb{R}^{2}\times(T^{2})_{34}$ to ×$\mathbb{R}\times(S^{1})_{2}\times(T^{2})_{34}$ with the size $L_{2}$, and we set $L=L_{3,4}\ll L_{2}\ll\Lambda^{-1}$. We then take the Hamiltonian viewpoint with regarding $\mathbb{R}$ as the time direction for the quantization Tanizaki:2022ngt ; Yamazaki:2017ulc ; Cox:2021vsa . As $L=L_{3,4}$ is the shortest length scale, we first minimize along those directions, and the ’t Hooft twist $n_{34}=1$ sets $P_{3}=S$, $P_{4}=C$ as we have already seen. Next, $P_{2}$ must commute with both $P_{3},P_{4}$ to minimize the classical action along the $2$-$3$ and $2$-$4$ directions, and then we obtain $N$ classical vacua labeled by $P_{2}=\mathrm{e}^{\frac{2\pi\mathrm{i}m}{N}}1_{N\times N}$ with $m=0,1,\ldots,N-1$. Let us call this state |$\ket{m}$:
@@
  • It is quite counter-intuitive that the Wilson loop in $2$d obeys the perimeter law whereas the vacuum structure is smoothly connected to the $4$d/$3$d confinement vacua. What is the microscopic dynamics realizing the switching555 The adiabatic continuity holds between small $L_{3}$ and large $L_{3}$ regimes, as shown in Section 3 by explicitly demonstrating that local observables and vacuum structure are kept under this compactification. However, this might sound to have a conflict with the switching of the Wilson loop behaviors from the area law to the perimeter law, which happens instantaneously upon the $\left(S^{1}\right)_{3}$ compactification regardless of how large $L_{3}$ is. Its consistency with the adiabatic continuity is ensured since this switching occurs for asymptotically large Wilson loops that are much larger than the compactification size $L_{3}$. from the 3d area law to the 2d perimeter law?
@@ 4.1.2 Wilson loop in ${\mathbb{R}}^{2}\times S^{1}$: perimeter law from double string picture
where $T_{\mathrm{kink}}$ is the kink tension, and $T_{\mathrm{conf}}$ is the string tension. If the size of $C$ is not so large compared with $L_{3}$, the standard area law appears. When the size of $C$ becomes large enough compared to $L_{3}$, the domain wall extended along $(S^{1})_{3}$ dominates and we obtain the perimeter law. The switching scale between these behaviors can be estimated as $T_{\mathrm{conf}}\,\mathrm{Area}(C_{*})=T_{\mathrm{kink}}\,L_{3}\,\mathrm{Length}(C_{*})$; when we use the double-string scenario with $T_{\mathrm{conf}}\approx 2T_{\mathrm{kink}}$, the switching size becomes $L_{*}\approx 2L_{3}$ for a square loop of size $L\times L$. This is the microscopic explanation for why the Wilson loops in the $2$d center-vortex semiclassics shows the perimeter law and generates the spontaneously-broken chiral transformation.
@@ 6 Summary and Discussion
On the one hand, we established the adiabatic continuity between the monopole and center-vortex semiclassical theories from $3$d to $2$d, with its explicit demonstration for dynamics, local observables, and vacuum structures. On the other hand, we also found the switching of the Wilson loop behaviors from the $3$d area law to $2$d perimeter law for finite $L_{3}$ as mentioned above. At the first sight, these two observations may seem to have a conflict. However, they are actually consistent because the switching from the $3$d area law to the $2$d perimeter law occurs only for asymptotically large Wilson loops, which are larger than the compactification size $L_{3}$. This is justified by estimating the switching size as $L_{*}\approx\frac{T_{\mathrm{conf}}}{T_{\mathrm{kink}}}L_{3}$.